Nay614:

the number -12,-7,-2,3,... follow a pattern. What are the next three numbers in this pattern?

jhonyy9:

what is the difference between these numbers ?

jhonyy9:

so how many you need adding to -12 to get -7 ?

Nay614:

the options are 8,14,17 5,10,15 2,8,13 8,13,18

jhonyy9:

ok but first pls answer my above question ?

jhonyy9:

-12 -(-7) = ?

jhonyy9:

or -12 +7 = ?

Nay614:

-19

jhonyy9:

how you get from -2 this next number of 3 ?

jhonyy9:

no -12 -7 = -19

Nay614:

ik

jhonyy9:

but there is -12 +7 = ?

Nay614:

its -19

Nay614:

thats what i got

jhonyy9:

how i ve wrote above -19 = -12 -7 but there is -12 -(-7) = -12 +7 = ?

jhonyy9:

or how you make from 3 this -2 3 + ? = -2

Nay614:

-5

jhonyy9:

perfect but from -12 how you get -7 ?

Timmyspu:

@jhonyy9 is right all the time about math so whatever he says he is usually right.

Nay614:

idk

Nay614:

and okkkk

jhonyy9:

why you say idk ? bc. the you ve got it when you wrote to 3 you need adding -5 t get -2 so this mean that the difference between the terms of this number line is 5

jhonyy9:

ok ?

jhonyy9:

so and now how you get the next term ?

Nay614:

ok i dont understand bc im butt at math

jhonyy9:

ok look pls -12 +5 = -7 -7 +5 = -2 -2 +5 = 3 3 +5 = ?

Nay614:

8

jhonyy9:

perfect the next term

Nay614:

?/

jhonyy9:

always add to the last 5

jhonyy9:

so the last was 8 and the next will be ... what ?

jhonyy9:

-2 +5 = 3 3 +5 = 8 ?

jhonyy9:

can somebody explain it more understandably pls ?